Jose invested $5,000.00. He invested part of it in a bond fund paying 5%, and the rest in a money market fund paying 4%. If the total interest earned for one year was $220.00, how much money did Jose invest in the money market fund?
Thank you,
lyra

Answer by ikdeep(226)

Note - here "*" is a mulitipication sign and "/" division sign

total money which Jose invested = $5,000.00
Let money invested in bond fund = $x..........equation no, (1)
Therefore, Money invested in market fund = ($5,000.00 - x).....Equation no,(2)
(because jose invested money in market fund which is left after investing in bond fund and we are given that he has totally invested $5,000.00 in both)
Now,total interest recieved by him = $220
Therefore according to question,we get the following equation:-
$220 = 5% of bond fund + 4% of market fund
220 = 5% of x + 4% of ($5,000.00 - x)
220 = 5/100 of x + 4/100 of (5,000.00 - x) (here 5/100 and 4/100 means that 5 and 4 are been divided by 100)

Since "of" is converted in Mutiply sign "*" and after solving brackets,,We get...
220 = 5/100 * x + 4/100 * 5,000.00 - 4/100 * x

now, after mutipication, we get ...
220 = 5/100 x + 20,000.00/100 - 4/100 x

Now, after taking L.C.M. we get
220 = (5x + 20,000.00 - 4x)/100

Now, after cross mutipication,we get...

22,000.00 = 5x + 20,000.00 - 4x
and 22,000.00 - 20,000.00 = 5x - 4x
2,000.00 = x
Now, put the value of x in equation no, (1)....we get
Money invested in Bond fund = $x
= $2,000.00
ANd Money invested in market fund = $5,000.00 - x
= $5,000.00 - 2,000.00
= $3,000.00
(Moreover if you want to check your answer. Take 5% of bond fund and 4% of market, the sum of both would have to be equal to $220 which is given total money recieved after investing in both the funds)